\begin{thebibliography}{00}

\bibitem{BD97} A. Bhatwadekar and A. Dutta,  {\em Kernels of locally nilpotent $R$-derivations on $R[X,Y]$.}  
Trans. Amer. Math. Soc. {\bf 349} (1997),  3003--3019, 
  

\bibitem{FM} D. Finston and S. Maubach, {\em Constructing (almost) rigid rings and a UFD having infinitely generated Derksen and Makar-Limanov invariant.}
Preprint,  2007

\bibitem{DF}
D. Daigle, G. Freudenburg, {\em A counterexample to Hilbert's fourteenth problem in dimension $5$.}  
J. Algebra  {\bf 221}  (1999),  no. 2, 528--535. 

\bibitem{R++} R. V. Gurjar, K. Masuda, M. Miyanishi, P. Russell, {\em Affine lines on affine surfaces and the Makar-Limanov invariant}, to appear in Cont.J.Math.

\bibitem{Mau} S. Maubach, {\em The linearisation conjecture and other problems over nonreduced rings.}  Comm. Algebra  {\bf 30}  (2002),   1693--1704. 

\bibitem{Kur1} S. Kuroda, {\em A counterexample to the fourteenth problem of Hilbert in dimension four.}  J. Algebra  {\bf 279}  (2004),   126--134. 

\bibitem{Kur2} S. Kuroda, {\em A counterexample to the fourteenth problem of Hilbert in dimension three.}  Michigan Math. J.  {\bf 53}  (2005),  123--132. 

%\bibitem{DF} J.K.Deveney, D.R.Finston, {\em $G_a$-actions on $\C^3$ and $\C^7$}, Comm.Alg 22 (1994), 37-41

%\bibitem{Essenboek} A.van den Essen, {\em Polynomial Automorphisms and the Jacobian Conjecture,} volume 190 of {\em in Progress in Mahtematics}, Birkh\"auser (2000)

%\bibitem{HY01}
%Pei-Chu Hu and Chung-Chun Yang, {\em Notes on a generalized abc-conjecture over function fields},
%Annales math\'ematiques Blaise Pascal, tome 8, no.1 (2001), 61-71


%\bibitem {FM06} D.Finston, S.Maubach, {\em The Automorphism Group of Certain Factorial Threefolds and a Cancellation Problem}, to appear in Isr.J.of Math.

%\bibitem {lenny}L. Makar-Limanov: \emph{\ Again }$x+x^{2}y+z^{2}+t^{3}%
%.$ Contemporary Mathematics 369, 177-182. \ American Mathematical Society,
%Providence (2005)

%\bibitem{scriptieIk} S.Maubach, {Hilbert's 14th problem and related topics,} master's thesis, University of Nijmegen (1998).


%\bibitem {S68}P. Samuel: \emph{Unique factorization}. Amer. Math. Monthly
%75 (1968) 945-952.




% \bibitem{label}
% Text of bibliographic item

%\bibitem{AM}
%S.S. Abhyankar, T.T. Moh, Embeddings of the line in the plane,
% J. Reine Angew. Math. 276 (1975) 148--166.

%\bibitem{BBvdE}
%J. Berson, J.W. Bikker, A. van~den Essen, Adapting coordinates, J.
%Pure Appl. Algebra 184 (2003) 165--174.

%\bibitem{BCMcC}
%J.W. Brewer, D.L. Costa, K. McCrimmon, Seminormality and root
%closure
% in polynomial rings and algebraic
%  curves, J. Algebra 58 (1979) 217--226.

%\bibitem{BD}
%S.M. Bhatwadekar, A.K. Dutta, On residual variables and stably
% polynomial algebras, Comm. Algebra 21 (1993) 635--645.

%\bibitem{Bh}
%S.M. Bhatwadekar, Generalized epimorphism theorem, Proc. Indian
%Acad.
 %Sci. Math. Sci. 98 (1988) 109--116.

%\bibitem{DY}
%V. Drensky, J.-T. Yu, Tame and wild coordinates of {$K[z][x,y]$},
% Trans. Amer. Math. Soc. 353 (2001) 519--537.

%\bibitem{EV}
%E. Edo, S. V\'en\'ereau, Length 2 variables of ${A}[x,y]$ and
%transfer, Ann. Polon. Math. 76 (2001) 67--76.

%\bibitem{Ha}
%E. Hamann, On the {$R$}-invariance of {$R[x]$}, J. Algebra 35 (1975)
%1--16.

%\bibitem{N2}
%M. Nagata, A theorem of {G}utwirth, J. Math. Kyoto Univ. 11 (1971)
%149--154.

%\bibitem{Na}
%M. Nagata, On automorphism group of $k[x,\,y]$, in: Kyoto University
% Lectures in Mathematics, Vol. 5, Kyoto University,
% Kinokuniya, Tokyo, 1972.

%\bibitem{RS}
%P. Russell, A. Sathaye, On finding and cancelling variables in
% $k[{x},\,{Y},\,{Z}]$, J. Algebra 57 (1979) 151--166.

%\bibitem{Ru}
%P. Russell, Simple birational extensions of two dimensional affine
% rational domains, Compositio Math. 33 (1976) 197--208  .

%\bibitem{Su}
%M. Suzuki, Propri\'et\'es topologiques des polyn\^omes de deux
%variables
%  complexes et automorphismes alg\'ebriques de l'espace
%${{\mathbb C}}\sp{2}$, J. Math. Soc. Japan 26 (1974) 241--257.

%\bibitem{KZ04}
%S. Kaliman,M. Zaidenberg,  V\'en\'ereau polynomials and related
%fiber bundles, J. Pure Appl. Algebra 192, No.1-3, 275-286 (2004).

%\bibitem{KVZ04}
%S. Kaliman, M. Zaidenberg, S. V\'en\'ereau, Simple birational
%extensions of the polynomial algebra $\C{[3]}$, Trans. Amer. Math.
%Soc, 356, No.2, 509-555 (2004).

%\bibitem{Sw}
%R.G. Swan, On seminormality, J. Algebra 67 (1980) 210--229.

%\bibitem{vdE}
%A. van den Essen, Polynomial automorphisms and the {J}acobian
%conjecture, Birkh\"auser Verlag, Basel, 2000.

%\bibitem{vdEvR}
%A. van~den Essen, P. van Rossum, Coordinates in two variables over a
 %{$\Bbb Q$}-algebra, Trans. Amer. Math. Soc. 356
% (2004) 1691--1703 (electronic).

%\bibitem{Ve02}
%S. V\'en\'ereau, Is $y+x[xz+y(yu+z^2)]$ an $x$-variable of
%$\C[x][y,z,u]?$, Problem session, Workshop on group action on
%rational varieties, (2002), 6p.
%\bibitem{Ve04}
%S. V\'en\'ereau, New bad lines in $R[x,y]$ and optimization of the
%Epimorphism Theorem, submitted (preprint available at
%http://www.math.unibas.ch/  under Aktivitaeten, Publikationen),
%(2004) 24 p.

%\bibitem{mth}
%S. V\'en\'ereau, Automorphismes et variables de l'anneau de
%polyn\^omes
%  $A[y_1,...,y_n]$, Ph.D. Thesis, Universit\'e Grenoble I, Institut
% Fourier, 2001.

% notes:
% \bibitem{label} \note

% subbibitems:
% \begin{subbibitems}{label}
% \bibitem{label1}
% \bibitem{label2}
% If there is a note, it should come last:
% \bibitem{label3} \note
% \end{subbibitems}

%\bibitem{}

\end{thebibliography}

\documentclass[12pt]{article}

\newlength{\stefan}
\usepackage{amsmath, amsthm, amsfonts, makeidx}
\usepackage{multind}
\DeclareMathSymbol{\subsetneq}{\mathord}{AMSb}{"26}

\newtheorem{lemma}{Lemma}[section]

\newtheorem{proposition}[lemma]{Proposition}

\theoremstyle{definition}

\newcommand{\mb}{\mathbb}

\newcommand{\C}{\mb{C}}

\newcommand{\N}{\mb{N}}

\title{Infinitely generated Derksen and Makar-Limanov invariant\footnote{AMS classification 13N15,13A50,14R20}}
\author{Stefan Maubach\footnote{Funded by a Veni-grant of the council for the
physical sciences, Netherlands Organisation for scientific research (NWO)}}

\begin{document}

\maketitle

\begin{abstract}
In this paper, we give an example of a finitely generated 3-dimensional $\C$-algebra which  has infinitely generated Derksen invariant as well as Makar-Limaonv invariant.
\end{abstract}

\section{Introduction and tools}

The Derksen invariant and Makar-Limanov invariant are useful tools to recognize if two varieties or rings are not isomorphic. 
Both invariants use locally nilpotent derivations: if $A$ is a commutative $k$-algebra (where $k$ is a field of characteristic zero), 
then $D$ is a derivation if $D$ is $k$-linear and satisfies the Leibniz rule: $D(ab)=aD(b)+bD(a)$. A derivation is locally nilpotent if 
for each $a\in A$ we can find some $n\in \N$ such that $D^n(a)=0$. The kernel of a derivation, denoted by $A^D$, is the set of all elements that are mapped 
to zero under the derivation $D$. The Makar-Limanov invariant is defined
as the intersection of all kernels of locally nilpotent derivations, while the Derksen invariant is defined 
as the smallest algebra containing the kernels of all nonzero locally nilpotent derivations.

In the paper \cite{R++} the question was posed if the Derksen invariant could be infinitely generated.
In this paper we give an example of an infinitely generated Derksen invariant of a finitely generated $\C$-algebra.
It will be at the same time an example of an infinitely generated Makar-Limanov invariant, as in this example, the Derksen invariant is equal to the Makar-Limanov invariant.
By now, there are many examples of cases of ``nice'' subrings that are not finitely generated \cite{BD97,DF,Mau,Kur1,Kur2}. In regard of this, the author would like to remark
that it will pay off to consider theorems as general as possible (with respect to not restricting to finitely generated algebras).

{\sl Notations:} If $R$ is a ring, then $R^{[n]}$ denotes the polynomial ring in $n$ variables over $R$.
We will use the letter $k$ for a field of characteristic zero, and $K$ for its algebraic closure.
Denote by $\partial_x$ the derivative with respect to $x$. By $LND(A)$ we will denote the set of all locally nilpotent derivations on a ring $A$. 

The technique used for constructing this example, is based on the following general  idea: grab a locally nilpotent derivation $D$ on a polynomial ring $K^{[n]}$ having the required
properties (non-finitely generated kernel).
Then construct an equation $f$ (or several equations $f_i$) which forces $A:=K^{[n]}/(f)$ (or $A:=K^{[n]}/(f_1,\ldots,f_m)$) 
to be a ring that has only one locally nilpotent derivation (up to multiplication with an element of $A$), namely $D \mod (f_1,\ldots,f_m)$. 
Hopefully it still has the required properties (infinitely generated kernel).

Well-known facts that we need  are the following:

\begin{lemma}\label{extra}
Let $D\in LND(A)$ where $A$ is a domain. \\
(1) Then $D(A^*)=0$.\\
(2) If $D(ab)=0$ with $a,b$ both nonzero, then $D(a)=D(b)=0$.\\
(3) If $fD\in LND(A)$ where $f\not = 0$ then $D(f)=0$ and $D\in LND(A)$.
\end{lemma}

\section{The  example}

This example is inspired by the example of Bhatwadekar and Dutta in \cite{BD97}.
We will write small letters for capital letters modulo a relation: for example, $a$ below is defined as $A+(A^3-B^2)$. 

Define $R:=\C[A,B]/(A^3-B^2)=\C[a,b]\cong\C[T^2,T^3]\subseteq \C[T]$.
Define $S:=R[X,Y,Z]/(Z^2-a^2(aX+bY)^2-1)=R[X,Y][z]$.
We leave it to the reader to check that $S$ is a domain.
We will first try to find all locally nilpotent derivations on this ring.

\begin{lemma}\label{E1.L1}
Let $D\in LND(S)$. Then $D(a)=D(b)=D(aX+bY)=D(z)=0$.
\end{lemma}

\begin{proof}
Since $(z-a^2X-abY)(z+a^2X+abY)=1$, we have by lemma \ref{extra} part 1 that $D(z-a^2X-abY)=D(z+a^2X+abY)=0$, and thus since $D$ is $\C$-linear,
$D(z)=D(a^2X+abY)=0$.
Because of lemma \ref{extra} part 2, we have $D(a)=D(aX+bY)=0$. Since $0=3a^2D(a)=D(a^3)=D(b^2)=2bD(b)$ we have $D(b)=0$.
\end{proof}

\begin{lemma}\label{E1.L2}
$LND(S)=S^D \cdot D$ where $D:= b\partial_X-a\partial_Y$.
\end{lemma}

\begin{proof}
Let $D\in LND(S)$. Then $aD(X)=-bD(Y)$ by lemma \ref{E1.L1}. Seeing $S$ as a subring of $B:=C\oplus C\bar{z}=C[Z]/(Z^2-T^8(X+TY)^2-1)$ where $C:=\C[T][X,Y]$, we write
$D(X)=f_0+\bar{z}f_1, D(Y)=g_0+\bar{z}g_1$ where $f_0,f_1, g_0,g_1\in C$. Since $D(X)=-TD(Y)$ we see that $T$ divides $f_0$ and $f_1$. But since both are in
$\C[T^2,T^3,X,Y]$ we know that even $T^2$ divides them. So, $D(X)=T^2h_0+T^2\bar{z}h_1$ where $h_i:=f_iT^{-2}$. But now $T(h_0+h_1\bar{z})=-g_0-g_1\bar{z}$, so $T$ divides the $g_i$.
Again, we have that even $T^2$ divides $g_i$, which then gives that $T$ divides $h_i$. In the end, $T^3$ divides $D(X)$ and $T^2$ divides $D(Y)$.
Write $D(X)=T^3f=T^3f_0+T^3zf_1, D(Y)=T^2g=T^2g_0+T^2zg_1$. Then $T^3f=D(X)=-TD(Y)=-T^3g$ so $f_0=-g_0,f_1=-g_1,$ and thus $f=-g$, and $D(X)=bf, D(Y)=af$. 
Since $D(z)=D(a)=D(b)=0$, $D=f( b\partial_X-a\partial_Y)$.
By lemma \ref{extra} we have that $f\in S^D$ and so we are done.
\end{proof}

\begin{proposition}\label{E1.P3}
Let $D:= T^3\partial_X-T^2\partial_Y$ on $S$ as before. Then  $S^D$ is not finitely generated as a $\C$-algebra.
\end{proposition}

\begin{proof}
Examining the natural extension of $D$ on $B:=\C[T][X,Y][z]$ it is easy to determine that $B^D=\C[T,z,X+TY]$.
Now $S^D=B^D\cap S$, as can be easily checked. Defining $P:=X+TY$, we are done if we show that $\C[T,z,P]\cap\C[T^2,T^3,z,X,Y]$ is not finitely generated.
We will do this by writing elements in a unique way in a representant system. \\
{\sc claim:} If $F\in S^D\backslash \C[T,z]$, then $F\in (T^2,T^3)S$. \\
{\sc proof of claim:} $S^D\mod(T^2,T^3)=\C[T,X+TY,z]/(T^2,z^2-1) \cap \C[z,X,Y]/(z^2-1)$. 
If $F\in S^D$ is nonzero, then 
\[ F\mod(T^2)=\sum_{i=0}^n f_i(\bar{T},\bar{z})(X+\bar{T}Y)^i =\sum_{i=0}^n f_i(\bar{T},\bar{z})(X^i+i\bar{T}X^{i-1}Y) \]
where $f_n\not = 0$. But since $F\mod(T^2,T^3)S \in \C[\bar{z},X,Y]$, this implies $n=0$. 
So $F\in \C[T,z]+(T^2,T^3)S$, which proves the claim.

If $S^D$ is finitely generated, then $S^D=\C[T^2,T^3,z,F_1,\ldots,F_n]$ where $F_i\in (T^2)B$ (by the claim).
Let $d$ be the maximum of the $X,Y$-degree of the $F_i$.
Take $T^2P^{d+1}$, which is in $S^D$.
Write $T^2P^{d+1}=f(F_1,\ldots,F_n)$ where $f$ has coefficients in $\C[T^2,T^3]\oplus
\C[T^2,T^3]z$. Computing modulo $(T^4)B$ (or $(T^4,T^5,T^6)S$) we see that
$T^2P^{d+1}=f_0+f_1F_1+\ldots+f_nF_n \mod{T^4}$ where $f_i\in \C[\bar{T}^2,\bar{T}^3]\oplus
\C[\bar{T}^2,\bar{T}^3]\bar{z}$, as $F_iF_j\in (T^4)B$ for all $i,j$. 
Now comparing the $X,Y$-degree from the right hand side (which is at most $d$) to the 
$X,Y$ degree  on the left hand side (which is $d+1$) we get a contradiction, showing that the assumption ``$S^D$ is finitely generated'' is wrong.
\end{proof}

{\bf Remark:}
The example in this paper is no UFD.
In the paper \cite{FM} an example of a $\C$-algebra UFD of dimension 6 is given, which has infinitely generated Derksen and Makar-Limanov invariant.

\noindent
{\sc Author's address}\\
\small
Stefan Maubach\\
Radboud University Nijmegen\\
\small Toernooiveld 1, The Netherlands\\ 
\small s.maubach@science.ru.nl

\end{document}